Determine the net radiant-heat exchange between two parallel oxidized iron plates when the distance between them is 2 ft. and each plate measures 10 by 10 ft. The surface of one is at 200 °F and the other is at `100°F.
1
Expert's answer
2021-10-27T07:02:26-0400

"T_1=200\u00b0F"

"T_2=100\u00b0F"

"Q_{net}=\frac{\sigma(T_1^4-T_2^4)}{\frac{1}{\epsilon_1}+\frac{1}{\epsilon_2}-1}"

"\epsilon_1=0.9"

"\epsilon_2=0.4"

"\sigma=5.67*10^{-8}"

"Q_{net}=\frac{5.67*10^{-8}(200^4-100^4)}{\frac{1}{0.9}+\frac{1}{0.4}-1}"

"=32.5723W\/ft^2"